At the beginning of a walk, Roberto and Juanita are 3.3 miles apart. If they leave at the same time and walk in the same direction, Roberto overtakes Juanita in 11 hours. If they walked towards each other, they meet in 1 hour. What are their speeds?
Answer by mananth(16949)

Distance = same= 3.3
Roberto speed =x
Jaunita speed =y
t=d/r
3.3 / ( x - y )= 11.00
11 ( x - y ) = 3.30
11 x -11 y = 3.3 ....................1
3.3 / ( x + y )= 1.00
1.00 ( x + y ) = 3.3
1.00 x + 1.00 y = 3.3 ...............2
Multiply (1) by 1.00
Multiply (2) by 11.00
we get
11 x + -11 y = 3.3
11 x + 11 y = 36.3
22 x = 39.6
/ 22
x = 1.8 mph Roberto speed

plug value of x in (1)
11 x -11 y = 3.3
19.8 -11 y = 3.3
-11 y = 3.3 -19.8
-11 y = -16.5
y = 1.5 mph Jaunita speed